# Prime Geodesic Theorem for Arithmetic Compact Surfaces

### v1

@@ 1. Introduction
where $[\gamma_{0}]$ runs through the set of primitive hyperbolic conjugacy classes in $\Gamma$, and $N([\gamma_{0}])=\alpha^{2}$ if $\gamma_{0}$ is conjugate to $\pm\begin{pmatrix}\alpha&\\ &\alpha^{-1}\end{pmatrix}$ for some $\alpha>1$. It is regarded as an analogue of the Riemann zeta function $\zeta(s)$. The norms $N([\gamma_{0}])$ are regarded as an analogue of prime numbers. In particular, their distribution satisfies the same asymptotic as the primes
@@
for some error term $E_{\Gamma}(x)$ smaller than $\mathrm{li}(x)$ as $x\to\infty$. An estimation like (1.1) is called a prime geodesic theorem for $\Gamma$. It is believed that the true order of the error term is $E_{\Gamma}(x)\ll_{\epsilon}x^{\frac{1}{2}+\epsilon}$. However the reason is different from the primes. Although $\mathrm{Z}_{\Gamma}(s)$ satisfies the Riemann hypothesis in certain cases, it is a meromorphic function of order $2$, hence has many more zeros than $\zeta(s)$. In fact the Riemann hypothesis for $\mathrm{Z}_{\Gamma}(s)$ implies only $E_{\Gamma}(x)\ll_{\epsilon}x^{\frac{3}{4}+\epsilon}$ (see [5] for example). Note that the counting function $\pi_{\Gamma}(x)$, like in the case of prime number theorem, is intimately related to another counting function of an analogue of the von-Mangoldt function
(1.2) $\Psi_{\Gamma}(x):=\sum_{\begin{subarray}{c}k\in\mathbb{Z}_{\geq 1}\\ [\gamma_{0}]:N([\gamma_{0}])^{k}\leq x\end{subarray}}\frac{\log N([\gamma_{0}])}{N([\gamma_{0}])^{\frac{k}{2}}-N([\gamma_{0}])^{-\frac{k}{2}}}=x+\widetilde{E}_{\Gamma}(x).$
In the special case $\Gamma={\rm PSL}_{2}(\mathbb{Z})$, many better bounds for $E_{\Gamma}(x)$ are available, e.g. $E_{\Gamma}(x)\ll_{\epsilon}x^{\frac{35}{48}+\epsilon}$ due to Iwaniec [7], $E_{\Gamma}(x)\ll_{\epsilon}x^{\frac{7}{10}+\epsilon}$ due to Luo–Rudnick–Sarnak [10], $E_{\Gamma}(x)\ll_{\epsilon}x^{\frac{71}{102}+\epsilon}$ due to Cai [2], and $E_{\Gamma}(x)\ll_{\epsilon}x^{\frac{25}{36}+\epsilon}$ due to Soundararajan–Young [11]. Note that the Luo–Rudnick–Sarnak bound is valid for any congruence subgroup of ${\rm PSL}_{2}(\mathbb{Z})$, and that the Soundararajan–Young bound is also valid for principal congruence subgroups of ${\rm PSL}_{2}(\mathbb{Z})$ by a previous work of the second author [3]. The generalization of these results to arbitrary co-compact lattices has various difficulties. In the case where $\Gamma=\Gamma_{\mathcal{D}}$ is the unit group of a maximal order of a division quaternion algebra $\mathcal{D}$ defined over $\mathbb{Q}$ and unramified at $\infty$, Koyama [9] extended Luo–Rudnick–Sarnk’s bound, and keeps the current record in this case.
@@
Koyama’s method exploits the Jacquet–Langlands correspondences, showing that the Laplacian eigenvalues on $\Gamma_{\mathcal{D}}\backslash\mathbb{H}$ are the same as those of new forms on $\Gamma_{0}(N)\backslash\mathbb{H}$ for some square-free integer $N$ determined by $\mathcal{D}$. This is an exploitation of the Jacquet–Langlands correspondences on the spectral sides. Its direct generalization to the (principal) congruence subgroups looks tricky. Our method exploits a basic tool in the trace formula proof of the Jacquet–Langlands correspondences [4], namely the matching of the orbital integrals. Concretely, we have two ingredients in our method. The first is a relation between the counting function $\Psi_{\Gamma}$ and the adelic stable orbital integrals for a class of lattices $\Gamma$ admitting a large normalizer $\widetilde{\Gamma}$. This is established in §3. The second is the matching of the adelic stable orbital integrals for a quaternion division algebra and the $2\times 2$ matrices. This matching is reduced to the relevant local ones, and is a variant of the local matching for ${\rm GL}_{2}$ and its inner form given in [4]. The major difficulty is to find the test functions on the $2\times 2$ matrices side. Our key observation is that, since the image of the local Jacquet–Langlands correspondences consists essentially only of supercuspidal representations, while the supercuspidal representations are constructed from chain orders, good candidates on the $2\times 2$ matrices side should be given by the principal congruence subgroups associated with these chain orders. We indeed find them in Proposition 2.4.
More interesting than the error term bound is the equation (4.5), which expresses the counting function $\Psi_{\Gamma_{\mathcal{D}}(\mathfrak{N})}(x)$ in terms of the counting functions $\Psi_{\Gamma}(x)$ for the relevant congruences subgroups $\Gamma$ on the $2\times 2$ matrices side. Further improvement of the error term bound depends only on those for such $\Gamma<{\rm PSL}_{2}(\mathbb{Z})$. In the case of Soundararajan–Young bound, such an improvement has non-trivial difficulties. We refer the reader to the computation for the Hecke congruence subgroups in [3]. It is also interesting to extend our result to non-principal congruence subgroups, which would require new input, too. On the other hand, a generalization of our method to co-compact quaternion lattices of ${\rm PSL}_{2}(\mathbb{C})$ should be easy.
@@ Acknowledgement
The authors would like to thank the organizers of the 2025 summer school “Algebra and Number Theory” held at the Chinese Academy of Sciences, during which event the paper is written. The second author would like to thank Jiandi Zou for inspiring discussion related to the chain orders.
@@ 2.1. Matching of Orbital Integrals
Let $\mathbf{F}$ be a $p$-adic field with normalized additive valuation $v$, valuation ring $\mathfrak{o}$, valuation ideal $\mathfrak{p}$ and a chosen uniformizer $\varpi$. Let $q:=\lvert\mathfrak{o}/\mathfrak{p}\rvert$. Up to isomorphism, there are two quaternion algebras over $\mathbf{F}$: ${\rm M}_{2}(\mathbf{F})$ and the unique division quaternion algebra $\mathcal{D}(\mathbf{F})$. Their reduced traces and reduced norms are denoted by ${\rm Tr}$, resp. ${\rm Tr}_{\mathcal{D}}$ and $\det$, resp. $\nu_{\mathcal{D}}$. Fix a uniformizer $\varpi_{\mathcal{D}}$ of $\mathcal{D}(\mathbf{F})$ with $\nu_{\mathcal{D}}(\varpi_{\mathcal{D}})=\varpi$. Their groups of invertible elements are ${\rm M}_{2}(\mathbf{F})^{\times}={\rm GL}_{2}(\mathbf{F})$ and $\mathcal{D}(\mathbf{F})^{\times}=\mathcal{D}^{\times}(\mathbf{F})$, respectively. We are particularly interested in their quotients by the center ${\rm PGL}_{2}(\mathbf{F})$ and ${\rm P}\mathcal{D}^{\times}(\mathbf{F}):=\mathcal{D}^{\times}(\mathbf{F})/\mathbf{F}^{\times}$, respectively.
Let $\mathcal{A}_{\mathcal{D}}(\mathbf{F})$, resp. $\mathcal{A}^{\lozenge}(\mathbf{F})$ be the set of irreducible unitary/admissible representations of ${\rm P}\mathcal{D}^{\times}(\mathbf{F})$, square-integrable representations of ${\rm PGL}_{2}(\mathbf{F})$. The following is (a special case of) the (local non-archimedean) Jacquet–Langlands correspondences for ${\rm PGL}_{2}$.
@@
In order to capture the above correspondences in the trace formulae, matching orbital integrals are introduced in [4, §8.B]. We reformulate a variant as follows, which can be viewed as the simultaneous matching for ${\rm GL}_{2}(\mathbf{F})$ and $\mathcal{D}^{\times}(\mathbf{F})$ with varying central characters.
@@
for irreducible admissible $\pi$ of ${\rm PGL}_{2}(\mathbf{F})$ that is neither square-integrable nor finite dimensional. Note that for a given $\varphi$ the existence of a/many matching $f$ is ensured by [4, Lemma (8.10)].
@@ 2.2. An Arithmetic Matching
Precisely, in ${\rm M}_{2}(\mathbf{F})$ we have two chain $\mathfrak{o}$-orders $\mathfrak{M}:={\rm M}_{2}(\mathfrak{o})$ and $\mathfrak{J}:=\begin{pmatrix}\mathfrak{o}&\mathfrak{o}\\ \mathfrak{p}&\mathfrak{o}\end{pmatrix}$. Up to conjugation they give all the two maximal compact subgroups modulo the center $\mathbf{F}^{\times}\mathfrak{M}^{\times}$ and $\mathbf{F}^{\times}\mathfrak{J}^{\times}$ of ${\rm PGL}_{2}(\mathbf{F})$. Let $\mathfrak{P}_{\mathfrak{M}}:=\varpi{\rm M}_{2}(\mathfrak{o})$, resp. $\mathfrak{P}_{\mathfrak{J}}=\begin{pmatrix}\mathfrak{p}&\mathfrak{o}\\ \mathfrak{p}&\mathfrak{p}\end{pmatrix}$ be the Jacobson radical of $\mathfrak{M}$, resp. $\mathfrak{J}$. Let $\mathbf{K}_{\mathfrak{M}}$, resp. $\mathbf{K}_{\mathfrak{J}}$ be the normalizer subgroup of $\mathfrak{M}$, resp. $\mathfrak{J}$ in ${\rm GL}_{2}(\mathbf{F})$. We have
@@
In $\mathcal{D}(\mathbf{F})$ we have the unique maximal $\mathfrak{o}$-order $\mathcal{O}_{\mathcal{D}}:=\left\{x\in\mathcal{D}(\mathbf{F})\ \middle|\ \nu_{\mathcal{D}}(x)\in\mathfrak{o}\right\}$. Then $\mathbf{F}^{\times}\mathcal{O}_{\mathcal{D}}^{\times}$ is the unique maximal compact subgroup modulo the center of ${\rm P}\mathcal{D}^{\times}(\mathbf{F})$. The Jacobson radical of $\mathcal{O}_{\mathcal{D}}$ is $\mathfrak{P}_{\mathcal{O}_{\mathcal{D}}}=\mathcal{P}_{\mathcal{D}}:=\varpi_{\mathcal{D}}\mathcal{O}_{\mathcal{D}}$. The normalizer subgroup of $\mathcal{O}_{\mathcal{D}}$ is $\mathcal{D}^{\times}(\mathbf{F})$.
@@
is satisfied only if the left hand side lies in $\mathbf{F}^{\times}\mathfrak{J}^{\times}$, since $\mathbf{B}(\mathbf{F})\cap\mathfrak{J}^{\times}\Pi=\emptyset$. Therefore for some $z\in\mathbf{F}^{\times}$
@@
$\mathfrak{M}^{\times}=\bigsqcup_{u\in\mathfrak{o}/\mathfrak{p}}w\begin{pmatrix}1&\\ u&1\end{pmatrix}\mathfrak{J}^{\times}\bigsqcup\mathfrak{J}^{\times}.$
@@
(2) Since $U_{\mathfrak{M}}^{n}\triangleleft\mathbf{K}_{\mathfrak{M}}$, the function $f_{n}$ is invariant by $\mathbf{K}_{\mathfrak{M}}$ by conjugation. We get by (1)
@@
We have the following formulae for $n\in\mathbb{Z}_{\geq 0}$

### v2

@@ 1. Introduction
where $[\gamma_{0}]$ runs through the set of primitive hyperbolic conjugacy classes in $\Gamma$, and $N([\gamma_{0}])=\alpha^{2}$ if $\gamma_{0}$ is conjugate to $\pm\begin{pmatrix}\alpha&\\ &\alpha^{-1}\end{pmatrix}$ for some $\alpha>1$. It is regarded as an analogue of the Riemann zeta function $\zeta(s)$. The conjugacy classes $[\gamma_{0}]$ are regarded as an analogue of prime numbers. In particular, their distribution satisfies the same asymptotic as the primes
@@
for some error term $E_{\Gamma}(x)$ smaller than $\mathrm{Li}(x)$ as $x\to\infty$. An estimation like (1.1) is called a prime geodesic theorem for $\Gamma$. It is believed that the true order of the error term is $E_{\Gamma}(x)\ll_{\epsilon}x^{\frac{1}{2}+\epsilon}$. However, the reason is different from the primes. Although $\mathrm{Z}_{\Gamma}(s)$ satisfies the Riemann hypothesis in certain cases, it is a meromorphic function of order $2$, hence has many more zeros than $\zeta(s)$. In fact the Riemann hypothesis for $\mathrm{Z}_{\Gamma}(s)$ implies only $E_{\Gamma}(x)\ll_{\epsilon}x^{\frac{3}{4}+\epsilon}$ (see [5] for example). Note that the counting function $\pi_{\Gamma}(x)$, like in the case of prime number theorem, is intimately related to another counting function of an analogue of the von-Mangoldt function
(1.2) $\Psi_{\Gamma}(x):=\sum_{k\in\mathbb{Z}_{\geq 1},[\gamma_{0}]:N([\gamma_{0}])^{k}\leq x}\frac{\log N([\gamma_{0}])}{1-N([\gamma_{0}])^{-k}}=x+\widetilde{E}_{\Gamma}(x).$
In the special case $\Gamma={\rm PSL}_{2}(\mathbb{Z})$, many better bounds for $E_{\Gamma}(x)$ are available, e.g. $E_{\Gamma}(x)\ll_{\epsilon}x^{\frac{35}{48}+\epsilon}$ due to Iwaniec [7], $E_{\Gamma}(x)\ll_{\epsilon}x^{\frac{7}{10}+\epsilon}$ due to Luo–Rudnick–Sarnak [10], $E_{\Gamma}(x)\ll_{\epsilon}x^{\frac{71}{102}+\epsilon}$ due to Cai [2], and $E_{\Gamma}(x)\ll_{\epsilon}x^{\frac{25}{36}+\epsilon}$ due to Soundararajan–Young [11]. Note that the Luo–Rudnick–Sarnak bound is valid for any congruence subgroup of ${\rm PSL}_{2}(\mathbb{Z})$, and that the Soundararajan–Young bound is also valid for principal congruence subgroups of ${\rm PSL}_{2}(\mathbb{Z})$ by a previous work of the second author [3]. The generalization of these results to arbitrary co-compact lattices has various difficulties. In the case where $\Gamma=\Gamma_{\mathcal{D}}$ is the unit group of a maximal order of a division quaternion algebra $\mathcal{D}$ defined over $\mathbb{Q}$ and unramified at $\infty$, Koyama [9] extended Luo–Rudnick–Sarnak’s bound, and keeps the current record in this case.
@@
Koyama’s method exploits the Jacquet–Langlands correspondences, showing that the Laplacian eigenvalues on $\Gamma_{\mathcal{D}}\backslash\mathbb{H}$ are the same as those of new forms on $\Gamma_{0}(N)\backslash\mathbb{H}$ for some square-free integer $N$ determined by $\mathcal{D}$. This is an exploitation of the Jacquet–Langlands correspondences on the spectral sides. Its direct generalization to the (principal) congruence subgroups looks tricky. Our method exploits a basic tool in the trace formula proof of the Jacquet–Langlands correspondences [4], namely the matching of the orbital integrals. Concretely, we have two ingredients in our method. The first is a relation between the counting function $\Psi_{\Gamma}$ and the adelic stable orbital integrals for a class of lattices $\Gamma$ admitting a large normalizer $\widetilde{\Gamma}$. This is established in §3. The second is the matching of the adelic stable orbital integrals for a quaternion division algebra and the $2\times 2$ matrices. This matching is reduced to the relevant local ones, and is a variant of the local matching for ${\rm GL}_{2}$ and its inner form given in [4]. The major difficulty is to find the test functions on the $2\times 2$ matrices side. Our key observation is the following. Since the image of the local Jacquet–Langlands correspondences consists essentially only of supercuspidal representations, while the supercuspidal representations are constructed from chain orders, good candidates on the $2\times 2$ matrices side should be given by the principal congruence subgroups associated with these chain orders. We indeed find them in Proposition 2.5.
More interesting than the error term bound is the equation (4.5), which expresses the counting function $\Psi_{\Gamma_{\mathcal{D}}(\mathfrak{N})}(x)$ in terms of the counting functions $\Psi_{\Gamma}(x)$ for the relevant congruences subgroups $\Gamma$ on the $2\times 2$ matrices side. Further improvement of the error term bound depends only on those for such $\Gamma<{\rm PSL}_{2}(\mathbb{Z})$. In the case of Soundararajan–Young bound, such an improvement has non-trivial difficulties. We refer the reader to the computation for the Hecke congruence subgroups in [3].
A generalization of our method to co-compact quaternion lattices of ${\rm PSL}_{2}(\mathbb{C})$ should be easy. It is also interesting to extend our result to non-principal congruence subgroups, which would require a new way of adelization without the crucial assumption (3.1) below. We formalize such possibility as a conjecture:
Conjecture 1.2.
For any congruence subgroup $\Gamma$ of ${\rm PSL}_{2}(\mathbb{R})$ defined by $\mathcal{D}$, there exist finitely many congruence subgroups $\Gamma_{i}<{\rm PSL}_{2}(\mathbb{Z})$ and $a_{i}\in\mathbb{Q}^{\times}$ so that
$\Psi_{\Gamma}(x)=\sum_{i}a_{i}\cdot\Psi_{\Gamma_{i}}(x),\quad\sum_{i}a_{i}=1.$
@@ Acknowledgement
The authors thank the referees for suggestions improving on the readability of the paper. They would like to thank the organizers of the 2025 summer school “Algebra and Number Theory” held at the Chinese Academy of Sciences, during which event the paper is written. Wu would like to thank Jiandi Zou for inspiring discussion related to the chain orders. Part of the work is done during Wu’s visit at the IASM of Zhejiang University. Wu would also like to thank Prof. Binyong Sun and IASM for hospitality and the support from the New Cornerstone Science Foundation.
@@ 2.1. Matching of Orbital Integrals
Let $\mathbf{F}$ be a $p$-adic field with normalized additive valuation $v$, valuation ring $\mathfrak{o}$, valuation ideal $\mathfrak{p}$ and a chosen uniformizer $\varpi$. Let $q:=\lvert\mathfrak{o}/\mathfrak{p}\rvert$. Up to isomorphism, there are two quaternion algebras over $\mathbf{F}$: ${\rm M}_{2}(\mathbf{F})$ and the unique division quaternion algebra $\mathcal{D}(\mathbf{F})$. Their reduced traces and reduced norms are denoted by ${\rm Tr}$, resp. ${\rm Tr}_{\mathcal{D}}$ and $\det$, resp. $\nu_{\mathcal{D}}$. Fix a uniformizer $\varpi_{\mathcal{D}}$ of $\mathcal{D}(\mathbf{F})$ with $\nu_{\mathcal{D}}(\varpi_{\mathcal{D}})=\varpi$. Their groups of invertible elements are ${\rm M}_{2}(\mathbf{F})^{\times}={\rm GL}_{2}(\mathbf{F})$ and $\mathcal{D}(\mathbf{F})^{\times}=\mathcal{D}^{\times}(\mathbf{F})$, respectively. We are particularly interested in their quotients by the center, ${\rm PGL}_{2}(\mathbf{F})$ and ${\rm P}\mathcal{D}^{\times}(\mathbf{F}):=\mathcal{D}^{\times}(\mathbf{F})/\mathbf{F}^{\times}$, respectively.
Let $\mathcal{A}_{\mathcal{D}}(\mathbf{F})$ and $\mathcal{A}^{\lozenge}(\mathbf{F})$ be the set of irreducible unitary representations of ${\rm P}\mathcal{D}^{\times}(\mathbf{F})$ and square-integrable representations of ${\rm PGL}_{2}(\mathbf{F})$, respectively. The following is (a special case of) the (local non-archimedean) Jacquet–Langlands correspondences for ${\rm PGL}_{2}$.
@@
In order to capture the above correspondences in the trace formulae, matching orbital integrals are introduced in [4, §8.B]. We reformulate a variant as follows.
@@
for irreducible admissible $\pi$ of ${\rm PGL}_{2}(\mathbf{F})$ that is neither square-integrable nor finite dimensional.
Remark 2.3.
If $f$ and $\varphi$ have a common central character, then for a given $\varphi$ the existence of a/many matching $f$ is ensured by [4, Lemma (8.10)]. For our version without central character, such existence is not known. The existence for our version obviously implies the one with central character.
@@ 2.2. An Arithmetic Matching
Precisely, in ${\rm M}_{2}(\mathbf{F})$ we have two chain $\mathfrak{o}$-orders $\mathfrak{M}:={\rm M}_{2}(\mathfrak{o})$ and $\mathfrak{J}:=\begin{pmatrix}\mathfrak{o}&\mathfrak{o}\\ \mathfrak{p}&\mathfrak{o}\end{pmatrix}$. Let $\mathfrak{P}_{\mathfrak{M}}:=\varpi{\rm M}_{2}(\mathfrak{o})$, resp. $\mathfrak{P}_{\mathfrak{J}}=\begin{pmatrix}\mathfrak{p}&\mathfrak{o}\\ \mathfrak{p}&\mathfrak{p}\end{pmatrix}$ be the Jacobson radical of $\mathfrak{M}$, resp. $\mathfrak{J}$. Let $\mathbf{K}_{\mathfrak{M}}$, resp. $\mathbf{K}_{\mathfrak{J}}$ be the normalizer subgroup of $\mathfrak{M}$, resp. $\mathfrak{J}$ in ${\rm GL}_{2}(\mathbf{F})$. Concretely we have
@@
Note that up to conjugation $\mathbf{K}_{\mathfrak{M}}$ and $\mathbf{K}_{\mathfrak{J}}$ are all the maximal compact subgroups of ${\rm PGL}_{2}(\mathbf{F})$. Note also that, since $\Pi^{2}=\varpi\mathbbm{1}_{2}$, $\mathfrak{o}^{\times}\mathbbm{1}_{2}<\mathfrak{J}^{\times}$ and $\Pi$ normalizes $\mathfrak{J}$ hence also $\mathfrak{J}^{\times}$ we have
$\mathbf{K}_{\mathfrak{J}}=\Pi^{2\mathbb{Z}}\mathfrak{J}^{\times}\bigsqcup\Pi^{2\mathbb{Z}+1}\mathfrak{J}^{\times}=\varpi^{\mathbb{Z}}\mathfrak{J}^{\times}\bigsqcup\varpi^{\mathbb{Z}}\mathfrak{J}^{\times}\Pi.$
In $\mathcal{D}(\mathbf{F})$ we have the unique maximal $\mathfrak{o}$-order $\mathcal{O}_{\mathcal{D}}:=\left\{x\in\mathcal{D}(\mathbf{F})\ \middle|\ \nu_{\mathcal{D}}(x)\in\mathfrak{o}\right\}$. Then $\mathbf{F}^{\times}\mathcal{O}_{\mathcal{D}}^{\times}$ is the unique maximal compact subgroup of ${\rm P}\mathcal{D}^{\times}(\mathbf{F})$. The Jacobson radical of $\mathcal{O}_{\mathcal{D}}$ is $\mathfrak{P}_{\mathcal{O}_{\mathcal{D}}}=\mathcal{P}_{\mathcal{D}}:=\varpi_{\mathcal{D}}\mathcal{O}_{\mathcal{D}}$. The normalizer subgroup of $\mathcal{O}_{\mathcal{D}}$ is $\mathcal{D}^{\times}(\mathbf{F})$.
@@
is satisfied only if the left hand side lies in $\mathbf{F}^{\times}\mathfrak{J}^{\times}$, since $\mathbf{B}(\mathbf{F})\cap\mathfrak{J}^{\times}\Pi=\emptyset$, where $\mathbf{B}(\mathbf{F})$ is the subgroup of upper triangular matrices of ${\rm GL}_{2}(\mathbf{F})$. Therefore for some $z\in\mathbf{F}^{\times}$
@@
$\mathfrak{M}^{\times}=\bigsqcup_{u\in\mathfrak{o}/\mathfrak{p}}w\begin{pmatrix}1&\\ u&1\end{pmatrix}\mathfrak{J}^{\times}\bigsqcup\mathfrak{J}^{\times},\quad w:=\begin{pmatrix}&1\\ 1&\end{pmatrix}$.
@@
(2) Since $U_{\mathfrak{M}}^{n}\triangleleft\mathbf{K}_{\mathfrak{M}}$, the function $f_{n}$ is invariant by conjugation by $\mathbf{K}_{\mathfrak{M}}$. We get by (1)
@@
Consider non-split $\mathbf{E}/\mathbf{F}$. We have the following formulae for $n\in\mathbb{Z}_{\geq 0}$